OpenStudy (anonymous):

Find f(g(x)) if f(x) = x - 4 and g(x) = x2.

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(g(x)) = f(x^2) = x^2 - 4}\)

Parth (parthkohli):

Factor that thing^

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - 2^2 = (x + 2)(x - 2) \text{ using difference of squares.}}\)

OpenStudy (anonymous):

is it x^2-4, x-4, x^2-2, or x^2+8x+16?

Parth (parthkohli):

As I said, \(x^2 - 4\)